### v1

@@ 3.1 Equivariant Graph Neural Networks
where and are embedding dimensions for node and edge, respectively. is represented by a multilayer perceptron, and is a random Fourier feature mapping [33] of the form,
where each parameter is sampled from the normal distribution . Embedding layers are followed by message passing layers defined by the following transformations
@@ 3.2 Training details
For the initial subsequence length , the loss function in DT-EGNN is defined as
@@
For subsequence lengths larger than one, the loss function includes only force terms
where the force term for each subsequence length is of the same form as in equation (11) and is equal to 50 for equal to one, and unity for any greater than one. Such scaling scheme proved to be crucial for the successful implementation of the method. This importance stems from the fact that during training process, initial structures are the only structures for which atomic positions exactly match those from the corresponding AIMD simulations.
@@
where is the total number of simulation steps and has the same form as in equation (11).
The loss function of the EGNN-MSE model has the following form
@@
All NNPs were trained on NVIDIA A100 GPUs in Python under version 3.12, Pytorch under version 2.4.0, and Pytorch Geometric under version 2.5.3. For the models trained with conventional methods (MACE, EGNN-MEPA, EGNN-MSE) we used 95%/5% splits resulting in 217,478 atomic configuration in the training set and 11,447 in the validation set. The learning rate, initially set to , was controlled by the Pytorch’s ReduceLRonPlateau scheduler. All NNPs were trained until the minimum learning rate of was reached except for the MACE NNP that was trained for 450 epochs. This constitutes nearly four times as many gradient updates than what is recommended in the official MACE documentation as a heuristic. Average epoch training times for the different NNPs employed in this study are summarized in Table 2. The reset value of the learning rate for DT-EGNN was set to . Training batch size was equal to 128 for the NNPs trained in a single-point fashion and 32 for DT-EGNN. Validation batch size was equal to 256 and test batch size was equal to one for all NNPs used in this work. All NNPs used Adam [34] as an optimizer.
@@
| EGNN-MSE | 675 |
4 Data availability
AIMD simulations and related informations used in the work have been deposited in the Figshare database under accession code https://doi.org/10.6084/m9.figshare.28498778.v1
5 Code availability
An open-source software implementation of DT-EGNN approach is available at https://github.com/IZugec/DTEGNN
@@ 6 Author contributions
7 Materials & Correspondence
Correspondance and requests for materials should be addressed to Ivan Žugec, Maite Alducin or J. Iñaki Juaristi.
Acknowledgements Financial support was provided by the Spanish MCINAEI10.13039501100011033, FEDER Una manera de hacer Europa (Grant No. PID2022-140163NB-I00), Gobierno Vasco-UPV/EHU (Project No. IT1569-22), and the Basque Government Education Departments’ IKUR program, also co-funded by the European NextGenerationEU action through the Spanish Plan de Recuperación, Transformación y Resiliencia (PRTR). Computer resources were provided by the Donostia International Physics Center (DIPC) Supercomputing Center.
@@ 7 Materials & Correspondence
Storing neighborhood structures at each step did not exceed our computational limits. However, applying the DT method to large datasets and systems with large amount of atoms could increase the computational cost. This is why we propose the following idea to alleviate the computational burden. If we assume that the changes in the neighborhood structure occur gradually along the dynamics, we can reuse the same neighborhood structure for multiple simulation steps. To examine the validity of this approximation for the system used in this work, we measure the rate of change of the neighborhood structure for atomic configurations offset by a different amount of simulation steps . As a measure of a difference between sets of neighbors we use the Jaccard similarity index
@@
Here we show how the choice of the loss function directly impacts the NNP performance during extended molecular dynamics simulations. The comparison involves three models: EGNN-MEPA, EGNN-MAE, and EGNN-MSE named after the loss functions used during training process. MEPA and MSE loss functions are already defined in the Method section of the main text and the mean absolute error (MAE) loss function is defined as

### v2

@@ 4.1 Equivariant Graph Neural Networks
where and are embedding dimensions for node and edge, respectively. is represented by a multilayer perceptron, and is a random Fourier feature mapping 33 of the form,
where each parameter is sampled from the normal distribution . Work by Tancik et al. 34 shows that such a mapping can overcome the spectral bias 35, 36 inherent to multilayer perceptrons. Embedding layers are followed by message passing layers defined by the following transformations
@@ 4.2 Training details
As shown schematically in Figure 1, the training of the potential energy surface starts by the traditional strategy of considering the different atomic structures in the dataset independently, without taking advantage of the dynamical information. In other words, each training data point has information on a single atomic structure and, in the context of our DT method, we denote them as subsequences of length one (). At this step, only the atomic positions, energies, and forces for each structure present in the dataset are needed and used for training. More precisely, for this initial subsequence length , the loss function in DT-EGNN is defined as
@@
After the convergence criteria is met for , the subsequent length is increased by one. In this case (), each data training point has information on two atomic structures, the initial one of the subsequence and the ensuing structure in the corresponding AIMD simulation. More precisely, the quantities extracted from the AIMD simulations and used during training are: the atomic positions and velocities of the first structure of the subsequence, the integration time step, and the forces for the two structures of each subsequence. Starting with the positions and velocities of the first structure of the subsequence, the model predicts the corresponding forces that, together with the integration time step, are used to generate the positions and velocities of the next step and predict the corresponding forces. The same procedure is used for larger subsequence lengths. In other words, only the time step, positions and velocities of the first structure of the subsequence are taken from the AIMD dataset. The positions and velocities of the rest of the structures in the subsequence are calculated using the velocity Verlet algorithm, and the forces of all the structures are predicted by the model. The training is performed by evaluating the accuracy of the model in predicting the forces of all the structures of the subsequence. More precisely, for subsequence lengths larger than one, the loss function is the following:
where the force term for each subsequence length is of the same form as in eq (11) and is equal to 50 for equal to one, and unity for any greater than one. Such scaling scheme proved to be crucial for the successful implementation of the method. This importance stems from the fact that during training process, initial structures are the only structures for which atomic positions exactly match those from the corresponding AIMD simulations. Note that, as described in Figure 1, the increase in the subsequence length is performed progressively. It is increased by one each time the force error convergence is met, until the predefined value is reached.
@@
where is the total number of simulation steps and has the same form as in eq (11).
The weighted mean square error loss function, used to train the MACE NNP, has the following form
where and are weights for the energy and force term respectively.
@@
All NNPs were trained on NVIDIA A100 GPUs in Python under version 3.12, Pytorch under version 2.4.0, and Pytorch Geometric under version 2.5.3. For the models trained with conventional methods (MACE, EGNN-MEPA) we used 95%/5% splits resulting in 217,478 atomic configuration in the training set and 11,447 in the validation set. The learning rate, initially set to , was controlled by the Pytorch’s ReduceLRonPlateau scheduler. All NNPs were trained until the minimum learning rate of was reached except for the MACE NNP that was trained for 450 epochs. This constitutes nearly four times as many gradient updates than what is recommended in the official MACE documentation as a heuristic. Average epoch training times for the different NNPs employed in this study are summarized in Table 2. The reset value of the learning rate for DT-EGNN was set to . Training batch size was equal to 128 for the NNPs trained in a single-point fashion and 32 for DT-EGNN. Validation batch size was equal to 256 and test batch size was equal to one for all NNPs used in this work. All NNPs used Adam 37 as an optimizer.
@@
5 Data and Software Availability
AIMD simulations and related informations used in the work have been deposited in the Figshare database under accession code https://doi.org/10.6084/m9.figshare.28498778.v1. An open-source software implementation of DT-EGNN approach is available at
https://github.com/IZugec/DTEGNN
@@ 6 Author contributions
7 Notes
The authors declare no competing financial interest.
Financial support was provided by the Spanish MCINAEI10.13039501100011033, FEDER Una manera de hacer Europa (Grant No. PID2022-140163NB-I00), Gobierno Vasco-UPV/EHU (Project No. IT1569-22), and the Basque Government Education Departments’ IKUR program, also co-funded by the European NextGenerationEU action through the Spanish Plan de Recuperación, Transformación y Resiliencia (PRTR). Computer resources were provided by the Donostia International Physics Center (DIPC) Supercomputing Center.
@@ 7 Notes
Storing neighborhood structures at each step did not exceed our computational limits. For this reason, this is the strategy that we followed in this work. The neighborhood structure was updated at every step according to the corresponding neighborhood structure in the AIMD simulations.
However, applying the DT method to large datasets and systems with large amount of atoms could increase the computational cost. In other words, for too large systems and too large datasets, storing the AIMD neighboorhood structures in order to update the neighboorhood structure at every step of the -loop could be prohibitive. This is why we propose the following idea to alleviate the computational burden. If we assume that the changes in the neighborhood structure occur gradually along the dynamics, we can reuse the same neighborhood structure for multiple simulation steps. To examine the validity of this approximation for the system used in this work, we measure the rate of change of the neighborhood structure for atomic configurations offset by a different amount of simulation steps . As a measure of a difference between sets of neighbors we use the Jaccard similarity index
@@
The results of Fig. 5 show that in the case of our system, instead of updating the atomic neighborhoods at every step, we could have opted for updating them after several steps. They also suggest that this could be the ideal strategy in case of too large systems for which updating at every step would be prohibitive. Note that, the main reason for this is that in well converged AIMD simulations the neigborhood structure varies smoothly. For example, the hydrogen atoms in the aforementioned AIMD simulations explore very different situations and neighborhoods (forming an isolated H2 molecule far from the Pd cluster, forming a H2 molecule interacting strongly with the Pd atoms, dissociated H atoms interacting with the Pd atoms). Still, the variation of the neighborhood structure for these H atoms along the dynamics is smooth enough to allow for updating neighborhood structure after several steps. This constitutes a strong indication of the validity of the proposed strategy for more general and larger systems of atoms.
@@
Here we show how the choice of the loss function directly impacts the NNP performance during extended molecular dynamics simulations. The comparison involves three models: EGNN-MEPA, EGNN-MAE, and EGNN-MSE named after the loss functions used during training process. Weighted MSE loss function, used to train MACE as well as EGNN-MSE, and MEPA are already defined in the Methods section of the main text. Mean absolute error (MAE) loss function is defined as
@@
Supplementary Note 3
Here we demonstrate the ability of DT-EGNN to capture a complex dynamical feature of the system that typically emerges only after thousands of integration steps. In Ref. 30 it was shown that hydrogen‐dissociation events are often accompanied by a structural rearrangement of the palladium cluster. More precisely, the dissociation of H2 in the Pd cluster can induce its transition from the original octahedral structure to an incomplete pentagonal bypiramid structure. It was shown that the structural transition basically consists in the drastic elongation of the Pd-Pd distance between two particular Pd neighbors in the equatorial plane of the octahedron. It is worth mentioning that this structural change is a relatively long time process that takes place 1-3 ps after H2 dissociation.
In order to test whether we can observe such events, we used the same initial conditions as the corresponding AIMD calculations and, using DT-EGNN as PES, propagated those initial conditions for 4 ps with the integration time step equal to 0.1 fs. Note that this requires performing 40 000 integration steps, to be compared with the 11 steps utilized in the longest training subsequences. The obtained results suggest that not only do we observe events featuring dissociation of the hydrogen molecule, but also reproduce the structural transition of the palladium cluster. Fig 9 shows the time evolution of the distance between the two relevant palladium atoms for both the DT-EGNN NNP and the AIMD reference. Although a sample of 100 trajectories is insufficient for statistically meaningfull discussion of dissociation probabilities (10 events for DT-EGNN vs. 7 for AIMD), it is noteworthy that three of these events (trajectories 6, 20 and 97) share identical initial conditions. All in all, these results show the good performance of the method on long time dynamics and its ability to describe complex dynamical phenomena.
Supplementary Note 4
Here we summarize the information on all DFT and computational settings that were employed in Ref. 30 to run the AIMD simulations, from which we obtained the dataset used in the present manuscript.
The calculations were based on spin-polarized DFT for a H2 molecule impinging on a Pd6 cluster anchored to a graphene vacancy (the system denoted as Pd6@G in the main text). The plane waves based VASP code 38, 39 was used to perform the calculations and exchange correlation effects were described using the generalized gradient approximation PW91 functional 40. The interaction of the explicitly considered valence electrons with the atomic cores was treated in the projector augmented wave (PAW) approximation 41, for which the PAW potentials supplied with the VASP package were used 42. The energy cutoff for the plane-wave basis set was 400 eV. The integration in the Brillouin zone was performed using a -centered 221 Monkhorst-Pack grid of special k-points 43. The first-order Methfessel-Paxton broadening scheme, with a 0.1 eV width, was used to consider fractional electronic-state occupancies. 44
The supercell consisted of a hexagonal 55 graphene layer in the lateral directions. The height of the supercell in the normal direction to the graphene layer was 14 Å. The substrate was relaxed until the forces on each Pd and C were below 0.02 eV/Å.
The ab initio molecular dynamic simulations used to generate the neural network potential energy surface in the present work were obtained for a H2 molecule impinging on Pd6@G under normal incidence conditions, with an initial substrate temperature of 300 K. The zero pressure limit was studied by considering one H2 per simulated trajectory. The initial translational and vibrational energies of the H2 molecule were E eV and Evib( 0, j0) = 0.27 eV, respectively. The center of mass of the molecules was initially located at a height 9 Åabove the graphene layer. The initial lateral position of the center of mass and the orientation of the molecule were randomly sampled. Before running the AIMD simulations, the Pd6@G substrate was equilibrated at 300 K for 7 ps using the Nosé thermostat 45. Sunsequently, the energy-conserving AIMD simulations of the H2 molecules impinging on the anchored cluster were performed, taking the initial positions and velocities of the C and Pd atoms randomly from the set of configurations generated during the thermalization at K. The number of calculated trajectories was 100, the integration time step 0.5 fs, and the total integration time was 4 ps.
Supplementary Note 5
Here, we show that our DT-EGNN NNP can perform molecular dynamics simulations lasting several orders of magnitude longer than those encountered during training, while maintaining the stability of the system. Additionally, we provide direct evidence for improved stability as the subsequence length, used during training process, increases.
Using DT-EGNN () as a NNP, 100 MD simulations with the integration time step of 0.5 fs and total time of 100 ps were performed. Taking the last snapshot of each of the 100 trajectories, we calculate distances between all atoms in the system and present them as a histogram. Furthermore, for comparison, we do the exact same thing for 100 simulations performed with AIMD. A comparison of normalized densities of interatomic distances between these two methods is shown on Fig. 10.
It is important to stress that the histogram for DT-EGNN is created for the atomic configurations at the end of 100 ps, while AIMD simulations last between 800 fs and 4 ps. Still, almost all of the peaks present in the AIMD structures are also present after 100 ps, providing compelling evidence that our DT-EGNN approach reliably maintains structural stability well beyond the subsequence lengths used during training.
Finally, we also performed 100 MD simulations using DT-EGNN at subsequence length equal to one () and compared the resulting distribution of interatomic distances at 100 ps to those obtained from AIMD simulations (see Fig. 11). Relative to DT-EGNN at , results for DT-EGNN at suggest that the discrete peaks of interatomic distances are less obvious for this distribution. Furthermore, significant portion of interatomic distances lay outside of the range of the AIMD distribution (very small as well as very big interatomic distances) suggesting much stronger deviations from the initial structure.